Speccy:

Help me solve! 10x - 25 = 45

NoodlesAndRiceYT:

i got this.

NoodlesAndRiceYT:

they dont call me human calculator for nun B)

NoodlesAndRiceYT:

ok so, what do you think happens first?

NoodlesAndRiceYT:

guys no i got this shush-

Speccy:

Subtract 25 from both sides!

NoodlesAndRiceYT:

no, you add not subtract, you do the oppisite for what -25 is

NoodlesAndRiceYT:

on both side

NoodlesAndRiceYT:

if it was +25, then you would subtract from both sides

NoodlesAndRiceYT:

get it?

Astrid1:

Ez pz I swear. 10x-25=45 is our equation. Youll want to ADD 25 to each side. 10x-25+25=45+25 Simplify. 10x=70 Divide each side by 10 10x/10 Obviously cancels out eachother. 70/10=7 x=7

NoodlesAndRiceYT:

now spec, what do you have now

NoodlesAndRiceYT:

OML ASTRID WHAT DID I JUST SAY

NoodlesAndRiceYT:

YALL REALLY JUST HAVE TO TAKE THE QUESTION AWAY FROM ME EH?

Speccy:

Ah yes, silly negatives. I didn't read that so keep going Soup.

NoodlesAndRiceYT:

I HAD IT ASTRID, DONT TAKE OTHER PEOPLES QUESTIONS, GOSH DANG

NoodlesAndRiceYT:

@speccy wrote:
Ah yes, silly negatives. I didn't read that so keep going Soup.
ok so, sense you added on both sides, what do you have now?

Speccy:

10x=70 divide both sides by 10 and x=7. Now that I have help I remembering.

NoodlesAndRiceYT:

amazing, you got it right!

NoodlesAndRiceYT:

im glad i helped you, if you need anymore help with it just ask meh